# Dynamic mapping from static labels: remote sensing dynamic sample generation with temporal-spectral embedding

### v1

@@ 3.5 Anomaly Attribution
To not only detect but also explain land surface dynamics in the dynamic samples, we further propose an anomaly interpretation method based on Gibbs sampling [Casella & George, 1992]. Specifically, after HTS-VAE identifies an anomalous time window, we aim to attribute the anomaly to specific spectral bands and temporal points, thereby pinpointing the sources of change within the dynamic sample. However, a major challenge arises: anomalous observations can contaminate the learned embeddings in HTS-VAE, leading to biased reconstructions even for the normal bands or time points. For example, an anomalous NDVI drop caused by cloud contamination may degrade the reconstruction of temporally adjacent or spectrally related bands such as NIR or Red, making it difficult to precisely localize the source of the change. To address this, we propose a Gibbs sampling-based anomaly attribution framework to refine the interpretation of detected anomalies by approximating what the “normal” reconstruction should be. The key idea is to iteratively update one spectral-temporal dimension at a time while conditioning on the fixed normal subset, allowing efficient and precise identification of the contribution of each dimension to the anomaly.
@@
3.6 Dynamic Sample Re-labeling
As shown in Fig. 3 (c), after we get the anomaly patterns and the interpretation results, we can relabel the temporal status of the samples to construct dynamic training samples. Here we consider an attribution significant if the reconstruction likelihood improves by more than a fixed threshold or falls within the top 10% of all attribution scores. Only significant anomalies are used to guide dynamic sample relabeling, ensuring that relabeling decisions are based on substantial spectral-temporal deviations rather than transient noise or minor perturbations. The core idea is to train a lightweight classifier to predict the class label of each time step within a sample, based on the spectral-temporal embeddings derived from the HTS-VAE. This enables us to transform a single static label (e.g., a known wetland class at time $t_{0}$) into a full sequence of dynamic labels $\{\hat{y}_{t}\}_{t=1}^{T}$ for the entire time series.
@@ 3.6 Dynamic Sample Re-labeling
Based on this, we successfully achieve ”Static to Dynamic”: Generating dynamic samples only using existing static annotations. The re-labeled sequences construct a dynamic sample pool, supporting downstream temporal modeling and dynamic land monitoring tasks such as dynamic land cover mapping and wetland mapping.
@@ 4.1 Study area
These regions are chosen based on the following considerations: First, most study areas exhibit a mixture of diverse land cover or wetland types, and the landscapes span different ecological and climatic zones across continents, providing a broad spectrum of surface dynamics. Second, high-resolution time-series imagery reveals distinct temporal evolution patterns across these regions. For example, SZ and CI exhibit rapid urban expansion within short periods, while NC remains relatively stable in terms of land cover. In the wetland mapping task, PL and MR show pronounced seasonal variations in wetland extent and type, whereas SD maintains a relatively stable wetland status throughout the year.
In summary, these six areas capture typical scenarios of land surface dynamics across various ecological and urban environments, featuring both homogeneous and heterogeneous landscapes. This diversity in land cover composition and surface change magnitude makes them well-suited for evaluating the land cover mapping and wetland mapping capabilities of TasGen, particularly its ability to automatically adapt to fine-grained temporal-spectral changes and generate temporally consistent classification results. These scenarios also provide a robust benchmark to assess the adaptability and accuracy of large-scale dynamic land cover classification and wetland mapping, especially when relying on automatically relabeled training data to track surface type transitions under complex and rapidly evolving environmental conditions.
@@ 5.3 Results of TasGen
Dynamic mapping. For the final dynamic mapping, TasGen achieves high overall accuracy (OA) across all regions, with scores above 89% in all cases. The Kappa coefficients also remain strong, validating not only the correctness of the predictions but also their statistical consistency with reference labels. Among the land cover regions, the CI region achieves the highest OA (93.78%) and Kappa (94.57%), likely due to clearer land use boundaries and lower intra-class variability. For wetland mapping, the PL region performs best OA (90.82%), showcasing TasGen’s ability to generalize to hydrologically dynamic and ecologically sensitive zones. From a vertical perspective, the mapping accuracy positively correlates with both F1-A and F1-S scores, reinforcing the importance of high-quality sample generation. TasGen’s ability to detect, correct, and relabel sample-level dynamics translates effectively into pixel-level classification outcomes. Fig. 9 shows the visualization results of dynamic mapping.
@@ 6 Discussion
6.1 Comparative Studies
To evaluate the effectiveness of TasGen in comparison with existing remote sensing time-series analysis methods, we conduct a comprehensive comparison against four widely-used baselines: CCDC [Zhu & Woodcock, 2014], LandTrendr [Kennedy et al., 2010], TSSCD [He et al., 2024], and static samples. CCDC models pixel-wise temporal trends using harmonic regression to continuously detect land surface changes and classify land cover types at each change point. LandTrendr segments spectral time series into linear pieces to detect long-term trends of disturbance and recovery. TSSCD leverages temporal dense annotations to model the temporal patterns using efficient deep learning networks, to perform semantic segmentation of land cover changes using long-range temporal context. Static samples use manually labeled static samples to train a classifier, assuming label consistency over time and lacking adaptability to dynamic land surface changes. Each method is applied to the same six regions and evaluated on the final dynamic mapping performance using Overall Accuracy (OA) and Kappa coefficient. Table 1 shows the comparison results.
Table 1: Comparison of mapping accuracy (OA / Kappa, in %) across six regions.
Task Region LandTrendr CCDC TSSCD Static TasGen (Ours)
Dynamic Land Cover Mapping SZ 86.13 / 85.20 88.35 / 87.02 94.42 / 91.45 86.05 / 83.76 92.14 / 91.97
CI 86.20 / 86.88 89.56 / 89.24 93.08 / 93.77 85.47 / 85.09 93.78 / 94.57
NC 88.94 / 86.31 87.32 / 88.56 92.41 / 91.98 86.29 / 85.23 91.54 / 89.59
Dynamic Wetland Mapping PL 83.43 / 81.77 84.69 / 84.23 88.55 / 86.32 82.18 / 81.35 90.82 / 91.84
MR 83.34 / 83.41 85.92 / 84.79 90.93 / 91.92 81.55 / 80.04 88.76 / 90.25
SD 84.18 / 84.33 85.35 / 85.68 90.47 / 89.19 81.09 / 82.13 90.69 / 89.36
TasGen achieves the best OA and Kappa in almost regions, with particularly large margins in complex wetland areas (PL, MR, SD), where static or change detection-based methods struggle to adapt to seasonal hydrological dynamics. This highlights the value of dynamic sample generation and anomaly-aware relabeling. Compared to CCDC and LandTrendr, which model pixel-level changes but ignore sample-level class transitions, TasGen leverages learned spectral-temporal embeddings and anomaly reasoning to reassign class labels precisely at the sample level. Although TSSCD achieves strong results, it relies on dense manual annotations and lacks sample-level interpretability. TasGen provides a more flexible, explainable, and annotation-efficient alternative, bridging change detection and supervised mapping. Static sample-based methods consistently underperform, reinforcing the need for dynamic modeling in rapidly evolving surface environments. Fig. 10 compares the change detection performance of TSSCD and CCDC in the same areas as Fig. 1 (b)–(d). Both methods exhibit notable misdetections and false alarms, especially under complex surface dynamics such as wetland transitions, compounded by dense observation frequencies (e.g., 4-day intervals). In contrast, although TasGen may also generate detection errors in the first stage, its subsequent sample relabeling module effectively mitigates these issues by the trained classifier. Therefore, TasGen can prioritize high recall during anomaly detection to ensure comprehensive change capture, while leveraging the second-stage relabeling to refine accuracy and reduce false positives. This comparison confirms that TasGen offers superior performance by trustable anomaly detection and accurate sample relabeling. Introducing dynamic samples as an interpretable and controllable intermediate variable increases the fault tolerance of anomaly detection and improves the accuracy of dynamic mapping.
6.2 Ablation Studies
@@
Anomaly detection. To evaluate the individual contribution of each module within TasGen, we conduct an ablation study across six geographically and ecologically diverse regions. Table 2 summarizes the anomaly detection F1-score (F1-A) under five model settings: the full TasGen architecture and four ablated variants where one key module is removed at a time.
@@
The full TasGen model consistently outperforms all ablated variants in every region, demonstrating the effectiveness of its integrated design. Removing the dual-dimension embedding structure leads to a significant performance drop. This module enables explicit compression of temporal and spectral dependencies into two disentangled embeddings, which are then fused hierarchically. Without this disentanglement, the model suffers from representational entanglement: temporal fluctuations may interfere with spectral pattern recognition, and vice versa. This makes it harder for the model to isolate true anomalies from natural variation in either domain. Disabling the HTS-VAE results in the largest drop in most regions. HTS-VAE introduces a structured generative process, where temporal context $\textbf{e}_{t}$ guides the learning of spectral representations $\textbf{e}_{s}$, ensuring that spectral anomalies are interpreted within their correct temporal phase. Without hierarchical conditioning, the model treats spectral and temporal embeddings independently, which leads to temporal misalignment and inaccurate anomaly detection. The sample curation strategy ensures that spectral embeddings are not learned from noisy or transient temporal patterns by first reconstructing a clean version of the time-series via Deconv1D layers. Removing this module means the model directly learns the patterns through the raw inputs, introducing risks of overfitting to clouds, sensor noise, or short-term fluctuations. This is especially problematic in wetland regions (e.g., PL, SD), where false positives are easily triggered by seasonal water-level shifts. The F1-A drops in these cases highlight the importance of pre-filtering as a noise-robust foundation for subsequent representation learning. Replacing the expressive Glow flow-based posterior with a standard Gaussian reduces model flexibility, resulting in moderate F1-A degradation. The Glow module expands the representational capacity beyond a simple Gaussian assumption, allowing the model to better approximate complex and multimodal latent distributions, which are common in real-world spatiotemporal data.
This ablation study confirms the necessity of each component in TasGen. Dual-dimensional embedding and HTS-VAE ensure structurally aligned temporal-spectral modeling, sample curation suppresses spurious signals, and Glow strengthens posterior modeling. Together, they form a synergistic framework for reliable and interpretable dynamic sample generation for remote sensing dynamic geographic mapping.
@@
Task Region w/o Dual-View w/o HTS-VAE w/o Curation w/o Glow Full (Ours)
@@
Dynamic sample relabeling. To assess the impact of different feature designs on dynamic sample relabeling, we conduct an ablation study by replacing the default feature $\mathbf{f}_{t}=\text{Fuse}(\mathbf{e}_{s_{t}},\mathbf{e}_{t_{t}})$ used by the classifier $\mathcal{C}$ with several alternative representations. This allows us to validate whether the proposed hierarchical temporal-spectral embeddings learned by HTS-VAE are indeed the most effective representation for label propagation across time steps.
@@ 6.3 Robustness Test
Table 4: Robustness test on anomaly detection F1-score (F1-A, in %) under (a) temporal frequency reduction and (b) random missing observations.
@@ 7 Conclusion
In this study, we propose TasGen, a novel two-stage framework for automatic dynamic sample generation in time-series dynamic geographic remote sensing mapping. By jointly modeling spectral and temporal dependencies through a hierarchical temporal-spectral variational autoencoder (HTS-VAE) and performing interpretable anomaly attribution using Gibbs sampling, TasGen achieves ”Static to Dynamic” with high interpretability and reliability: effectively transforming static annotations into temporally dynamic training samples for dynamic mapping. Extensive experiments across six geographically diverse regions and two mapping tasks, i.e., land cover and wetland mapping, demonstrate that TasGen consistently outperforms existing state-of-the-art methods in anomaly detection, sample relabeling, and final mapping accuracy. The method shows strong robustness to observation sparsity and missing data, while maintaining high adaptability in different surface dynamics. By reducing the dependence on continuous manual annotations and enhancing the temporal validity of training data, TasGen provides a scalable and accurate solution for dynamic remote sensing monitoring, offering new possibilities for large-scale land change analysis and environmental applications.

### v2

@@ 2.6 Anomaly attribution
To not only detect but also explain land surface dynamics in the dynamic samples, we further need to propose an anomaly interpretation method to attribute detected anomalies. Specifically, after HTS-VAE identifies an anomalous time window, we aim to attribute the anomaly to specific spectral bands and temporal points, thereby pinpointing the sources of change within the dynamic sample. However, a major challenge arises: anomalous observations can contaminate the learned embeddings in HTS-VAE, leading to biased reconstructions even for the normal bands or time points. For example, an anomalous NDVI drop caused by cloud contamination may degrade the reconstruction of temporally adjacent or spectrally related bands such as NIR or Red, making it difficult to precisely localize the source of the change. To attribute detected anomalies to specific spectral bands and time points, we require a method that can isolate individual spectral-temporal dimensions while preserving the context of the remaining data. Full-sample reconstruction may obscure localized signals, whereas dimension-wise analysis enables finer attribution. Gibbs sampling [Casella & George, 1992] offers an effective solution by iteratively refining one dimension at a time, conditioned on the rest. This allows us to approximate what each anomalous value would be under normal conditions, without disrupting the overall data structure. Based on this, we propose a Gibbs sampling-based anomaly attribution framework to refine the interpretation of detected anomalies by approximating what the “normal” reconstruction should be. The key idea is to iteratively update one spectral-temporal dimension at a time while conditioning on the fixed normal subset, allowing efficient and precise identification of the contribution of each dimension to the anomaly.
@@
2.7 Dynamic sample relabeling
As shown in Fig. 3 (c), after we get the anomaly patterns and the interpretation results, we can relabel the temporal status of the samples to construct dynamic training samples. Here we consider an attribution significant if the reconstruction likelihood improves by more than a fixed threshold or falls within the top 10% of all attribution scores. Only significant anomalies are used to guide dynamic sample relabeling, ensuring that relabeling decisions are based on substantial spectral-temporal deviations. We train a lightweight classifier to predict the class label of each time step within a sample, based on the spectral-temporal embeddings derived from the HTS-VAE. This enables us to transform a single static label (e.g., a known wetland class at time $t_{0}$) into a full sequence of dynamic labels $\{\hat{y}_{t}\}_{t=1}^{T}$ for the entire time series.
@@ 2.7 Dynamic sample relabeling
Based on this, we successfully achieve ”Static to Dynamic”: Generating dynamic samples only using existing static annotations. The re-labeled sequences construct a dynamic sample pool, supporting temporal modeling and dynamic land monitoring (i.e., dynamic land cover mapping and wetland mapping).
@@ 3.1 Study area
These regions are chosen based on the following considerations: First, most study areas exhibit a mixture of diverse land cover or wetland types, and the landscapes span different ecological and climatic zones across continents, providing a broad spectrum of surface dynamics. Besides, the areas of these regions are large enough ($12,062\ km^{2}$ of each) to capture significant spatial heterogeneity and representativeness, enabling robust evaluation of dynamic mapping methods at multiple scales. Second, there are distinct temporal evolution patterns across these regions. For example, SZ and CI exhibit rapid urban expansion within short periods, while NC remains relatively stable in terms of land cover. In the wetland mapping task, PL and MR show pronounced seasonal variations in wetland extent and type, whereas SD maintains a relatively stable wetland status throughout the year.
@@
In summary, these six areas capture typical scenarios of land surface dynamics across various ecological and urban environments, featuring both homogeneous and heterogeneous landscapes. This diversity in surface composition and surface change magnitude makes them well-suited for evaluating the land cover mapping and wetland mapping capabilities of TasGen, particularly its ability to automatically adapt to fine-grained temporal-spectral changes and generate temporally consistent classification results. These scenarios also provide a robust benchmark to assess the adaptability and accuracy of large-scale dynamic land cover classification and wetland mapping, especially when relying on automatically relabeled training data to track surface type transitions under complex and rapidly evolving environmental conditions.
@@ 4.2 Metrics
4.3 Results of TasGen
TasGen generates dynamic samples via two stages: anomaly detection and dynamic sample relabeling. After getting dynamic samples, dynamic mapping is conducted through a trained classifier. Therefore, we evaluate the performance of TasGen based on anomaly detection, dynamic sample relabeling, and dynamic mapping. We also visualize the performance of the anomaly attribution module in this section.
@@ 4.3 Results of TasGen
Dynamic mapping. For the final dynamic mapping, TasGen achieves high overall accuracy (OA) across all regions, with scores above 89% in all cases. The Kappa coefficients also remain strong, validating not only the correctness of the predictions but also their statistical consistency with reference labels. Among the land cover regions, the CI region achieves the highest OA (93.44%) and Kappa (0.925), likely due to clearer land use boundaries and lower intra-class variability. For wetland mapping, the PL region performs best OA (91.82%), showcasing TasGen’s ability to generalize to hydrologically dynamic and ecologically sensitive zones. From a vertical perspective, the mapping accuracy positively correlates with both F1-A and F1-S scores, reinforcing the importance of high-quality sample generation. TasGen’s ability to detect, correct, and relabel sample-level dynamics translates effectively into pixel-level classification outcomes. Fig. 9 shows the visualization results of dynamic mapping.
@@
4.4 Comparative studies
To evaluate the effectiveness of TasGen in comparison with existing remote sensing time-series analysis methods, we conduct a comprehensive comparison against four widely-used baselines: CCDC [Zhu & Woodcock, 2014], LandTrendr [Kennedy et al., 2010], TSSCD [He et al., 2024], and static samples. CCDC models pixel-wise temporal trends using harmonic regression to continuously detect land surface changes and classify land cover types at each change point. LandTrendr segments spectral time series into linear pieces to detect long-term trends of disturbance and recovery. TSSCD leverages temporal dense annotations to model the temporal patterns using efficient deep learning networks, to perform semantic segmentation of land cover changes using long-range temporal context. Static samples use manually labeled static samples to train a classifier, assuming label consistency over time and lacking adaptability to dynamic land surface changes. Each method is applied to the same six regions and evaluated on the final dynamic mapping performance using Overall Accuracy (OA) and Kappa coefficient. Table 1 and Table 2 show the comparison results.
Table 1: Comparison of mapping accuracy (OA in %) across six regions.
Task Region LandTrendr CCDC TSSCD Static TasGen (Ours)
Dynamic Land Cover Mapping SZ 85.91 ± 0.36 87.64 ± 0.37 91.93 ± 0.31 85.71 ± 0.46 92.07 ± 0.39
CI 86.01 ± 0.35 89.49 ± 0.51 92.83 ± 0.53 85.57 ± 0.37 93.44 ± 0.43
NC 88.74 ± 0.29 87.58 ± 0.23 92.17 ± 0.39 86.31 ± 0.36 91.11 ± 0.61
Dynamic Wetland Mapping PL 83.35 ± 0.62 84.12 ± 0.69 88.14 ± 0.61 82.19 ± 0.54 91.82 ± 0.70
MR 83.86 ± 0.45 85.46 ± 0.56 89.64 ± 0.64 81.43 ± 0.35 90.17 ± 0.79
SD 84.48 ± 0.45 85.64 ± 0.67 90.01 ± 0.58 80.82 ± 0.56 91.25 ± 0.54
Table 2: Comparison of mapping accuracy (Kappa) across six regions.
Task Region LandTrendr CCDC TSSCD Static TasGen (Ours)
Dynamic Land Cover Mapping SZ 85.01 ± 0.36 86.35 ± 0.21 90.14 ± 0.50 84.05 ± 0.36 90.54 ± 0.59
CI 85.61 ± 0.45 88.02 ± 0.37 92.01 ± 0.63 85.47 ± 0.27 92.52 ± 0.53
NC 87.82 ± 0.39 87.32 ± 0.43 91.46 ± 0.58 86.02 ± 0.36 91.01 ± 0.31
Dynamic Wetland Mapping PL 83.13 ± 0.32 83.69 ± 0.69 87.85 ± 0.65 82.08 ± 0.54 90.05 ± 0.63
MR 83.37 ± 0.41 85.22 ± 0.36 89.23 ± 0.53 80.79 ± 0.45 89.86 ± 0.60
SD 84.12 ± 0.25 85.21 ± 0.47 89.61 ± 0.54 80.16 ± 0.46 90.29 ± 0.64
TasGen achieves the best OA and Kappa in almost regions, with particularly large margins in complex wetland areas (PL, MR, SD), where static or change detection-based methods struggle to adapt to seasonal hydrological dynamics. This highlights the value of dynamic sample generation and anomaly-aware relabeling. Compared to CCDC and LandTrendr, which model pixel-level changes but ignore sample-level class transitions, TasGen leverages learned spectral-temporal embeddings and anomaly reasoning to reassign class labels precisely at the sample level. Although TSSCD achieves strong results, it relies on dense manual annotations and lacks sample-level interpretability. TasGen provides a more flexible, explainable, and annotation-efficient alternative, bridging change detection and supervised mapping. Static sample-based methods consistently underperform, reinforcing the need for dynamic modeling in rapidly evolving surface environments. Fig. 10 visualizes the dynamic mapping results of TasGen, TSSCD and CCDC. Fig. 11 compares the change detection performance of TSSCD and CCDC in the same areas as Fig. 7 (b)–(d). Both methods exhibit notable misdetections and false alarms, especially under complex surface dynamics such as wetland transitions, compounded by dense observation frequencies (e.g., 4-day intervals). In contrast, although TasGen may also generate detection errors in the first stage, its subsequent sample relabeling module effectively mitigates these issues by the trained classifier. Therefore, TasGen can prioritize high recall during anomaly detection to ensure comprehensive change capture, while leveraging the second-stage relabeling to refine accuracy and reduce false positives. This comparison confirms that TasGen offers superior performance by trustable anomaly detection and accurate sample relabeling. Introducing dynamic samples as an interpretable and controllable intermediate variable increases the fault tolerance of anomaly detection and improves the accuracy of dynamic mapping.
@@ 5 Discussion
5.1 Effectiveness of modules in anomaly detection
@@
To evaluate the individual contribution of each module within TasGen, we conduct an ablation study across six geographically and ecologically diverse regions. Table 3 summarizes the anomaly detection F1-score (F1-A) under five model settings: the full TasGen architecture and four ablated variants where one key module is removed at a time.
@@
Rationale for decoupling and coupling. Removing the dual-dimension embedding means we directly learn the temporal and spectral features jointly, without decoupling them in the first stage. Removing the HTS-VAE means we learn the temporal and spectral patterns separately, without recognizing them as a whole in the reconstruction process. We can see from Table 3 that removing the dual-dimension embedding structure leads to a significant performance drop. This module enables explicit compression of temporal and spectral dependencies into two disentangled embeddings, which are then fused hierarchically. Without this disentanglement, the model suffers from representational entanglement: temporal fluctuations may interfere with spectral pattern recognition, and vice versa. This makes it harder for the model to isolate true anomalies from natural variation in either domain. Disabling the HTS-VAE results in the largest drop in most regions. HTS-VAE introduces a structured generative process, where temporal context $\textbf{e}_{t}$ guides the learning of spectral representations $\textbf{e}_{s}$, ensuring that spectral anomalies are interpreted within their correct temporal phase. Without hierarchical conditioning, the model treats spectral and temporal embeddings independently, which leads to temporal misalignment and inaccurate anomaly detection.
@@
Task Region w/o Dual-dimension w/o HTS-VAE w/o Curation w/o Glow Full (Ours)
@@
Rationale for sample curation. The sample curation strategy ensures spectral embeddings are not learned from noisy temporal patterns by reconstructing a clean time-series via Deconv1D layers. Removing this module means the model directly learns the patterns through the raw inputs, introducing risks of overfitting to clouds, sensor noise, or short-term fluctuations. This is especially problematic in wetland regions (e.g., PL, SD), where false positives are easily triggered by seasonal water-level shifts. The F1-A drops in these cases highlight the importance of pre-filtering as a noise-robust foundation for subsequent representation learning.
Rationale for Glow. Replacing the Glow flow-based posterior with a standard Gaussian reduces model flexibility, resulting in moderate F1-A degradation. The Glow module expands the representational capacity beyond a simple Gaussian assumption, allowing the model to better approximate complex and multimodal latent distributions, which are common in real-world spatiotemporal data.
The full TasGen model consistently outperforms all ablated variants in every region, demonstrating the effectiveness of its integrated design. This ablation study confirms the necessity of each component in the anomaly detection stage of TasGen. Dual-dimensional embedding and HTS-VAE ensure structurally aligned temporal-spectral modeling, sample curation suppresses spurious signals, and Glow strengthens posterior modeling. Together, they form a synergistic framework for reliable anomaly detection for dynamic sample generation later.
5.2 Effectiveness of feature design in dynamic sample relabeling
To further assess the impact of different feature designs on dynamic sample relabeling, we conduct an ablation study by replacing the default feature $\mathbf{f}_{t}=\text{Fuse}(\mathbf{e}_{s_{t}},\mathbf{e}_{t_{t}})$ used by the classifier $\mathcal{C}$ with several alternative representations. This allows us to validate whether the proposed hierarchical temporal-spectral embeddings learned by HTS-VAE are indeed the most effective representation for label propagation across time steps.
@@ 5.3 Robustness test
Table 5: Robustness test on anomaly detection F1-score (F1-A, %) under reduced observation frequency (1/2 obs, 1/4 obs, 1/8 obs) and random missing observations (10%, 30%, 50%).
@@ 6 Conclusion
To enable scalable, interpretable, and accurate dynamic geographic mapping with minimal manual effort, in this study, we propose TasGen, a novel two-stage framework for automatic dynamic sample generation in time-series dynamic remote sensing mapping. By first decoupling and then jointly modeling spectral and temporal dependencies through a dual-dimension embedding and a hierarchical temporal-spectral variational autoencoder (HTS-VAE), with performing interpretable anomaly attribution using Gibbs sampling, and relabeling the dynamic samples via a trained classifier, TasGen achieves ”Static to Dynamic” with high interpretability and reliability: effectively transforming static annotations into temporally dynamic training samples for dynamic mapping with no more human intervention. Extensive experiments across six geographically diverse regions and two mapping tasks, i.e., land cover and wetland mapping, demonstrate that TasGen consistently achieves an average OA of 95.08%, 95.01%, and 91.64% in anomaly detection, dynamic sample generation, and dynamic mapping, respectively, outperforming existing state-of-the-art methods. Based on stable classification theory [Gong et al., 2024], generated dynamic samples can serve as reliable supervision for high-quality dynamic mapping, which is also proven in our experiments. Comprehensive ablation studies illustrate the effectiveness of modules and designs. The method shows strong robustness to observation sparsity and missing data, while maintaining high adaptability in different surface dynamics. By reducing the dependence on continuous manual annotations and enhancing the temporal validity of training data, TasGen provides a scalable and accurate solution for dynamic remote sensing monitoring, offering new possibilities for large-scale land change analysis and environmental applications.